OpenStudy (anonymous):

If f(x) = 12x + 7, what is f^–1(x) ?

OpenStudy (anonymous):

would it be f^-1 (12x + 7) ?

OpenStudy (saifoo.khan):

y = 12x + 7 Solve for "x".

OpenStudy (hoblos):

y=12x +7 switch x & y and solve for y

OpenStudy (hoblos):

x=12y +7 12y=x-7 y=(1/12)x - 7/12 so f^–1(x)= (1/12)x - 7/12